# GHP-MOFassemble: Diffusion modeling, high throughput screening, and molecular dynamics for rational discovery of novel metal-organic frameworks for carbon capture at scale

### v1

@@ 4.2 Linker generation and evaluation
0.995
0.798 0.789
4.3 MOF assembly
We generate new MOF structures by assembling three randomly selected DiffLinker-generated linkers with one of the three most frequent nodes in the hMOF dataset. Random sampling of linkers ensures that the selected linkers cover a large design space. For each node-linker-topology combination, we considered four levels of catenation (cat0, cat1, cat2, cat3). We generated a total of 120,000 MOFs as follows: we have four catenation levels and three node candidates. The random sampling of 10,000 linkers for each catenation level-node candidate pair generates 120,000 total MOFs of different catenation-node-linker combinations. We provide more details in Section 3.2.3 on how these new MOFs were assembled.
4.4 Regression model for MOF $\textrm{CO}_{2}$ capacity prediction
4.4.1 Pre-training on hMOF structures
We employ the modified version of the CGCNN model [18] as a predictive model of the CO2 capacity of MOF structures. We trained this model on 90% of the MOF structures in hMOF, along with their CO2 capacities at 0.1 bar, leaving 10% as a holdout set for testing. We trained the modified version of the CGCNN model three times to create an ensemble of models in order to mitigate biases in individual models. We summarize in Table 6 and Figure 12 the prediction errors of each model on the 10% test set and the distribution of standard deviations, respectively.
Looking next at the ensemble model, we show in the left panel of Figure 7 its predictions for the 10% test set. The ensemble model has a mean absolute error (MAE) of 0.093 mmol/g on the test set. Using CO2 capacity of 2 mmol/g as a threshold, we repurposed our predictive model as a classifier to categorize MOFs into low and high performers (with predicted CO2 capacities below and above the threshold, respectively) and high performers. Using this scheme, the confusion matrix for identifying low and high performers is shown in the right panel of Figure 7.
We conclude from the confusion matrix that the pre-trained model classifies both low and high performers with high accuracy, with 98.4% (13551 out of 13765) of test samples correctly classified. As the test set is heavily imbalanced, with many more low performers than high performers, we also calculated the balanced accuracy of classification [51] (the sum of true positive rate and true negative rate divided by 2), obtaining a value of 90.7%. The majority of the 214 misclassified samples lie close to the decision boundaries, as shown in the red and purple regions of the left panel of Figure 7. Therefore, we also conclude that the ensemble model is capable of differentiating low and high performers.
Figure 7: Left panel, predictive performance of the ensemble model on the 10% test set. Right panel, confusion matrix of the ensemble model in classifying low and high performers in the test set. The classification threshold is 2 mmol/g.
4.4.2 Predictions for newly generated MOF structures
When using the ensemble model to infer the CO2 capacities of newly generated MOF structures, we employ the average of the predictions made by the three independent models as the predicted capacity. In addition, we filter out structures for which the standard deviation of the predicted values obtained from the three models is larger than 0.2 mmol/g.
We treat the standard deviation of predictions from the three models as a measure of the uncertainty of model predictions, which we view as arising from the model’s difficulty in learning certain data points. We chose the threshold for the standard deviation filter of 0.2 mmol/g because it is sufficiently small (with 96% of data points below the threshold) that for low CO2 capacity predictions, the predicted values across the ensemble are similar. On the other hand, for high CO2 capacity predictions, it also ensures that outlier values (e.g., predictions made by the three models are vastly different) are filtered out, therefore minimizing the overall error.
We show in Figure 8 the empirical cumulative distribution functions of both the hMOF (dashed lines) and predicted (sold lines) CO2 capacities. We observe that there are more high-performing cat2 and cat3 MOFs among the generated structures than in the hMOF dataset, as indicated by lower cumulative density values at 2 mmol/g (black dashed line). On the other hand, there are fewer high-performing cat0 and cat1 MOFs, as indicated by higher cumulative density values at 2 mmol/g. This result indicates that within the generated MOF design space, there are more high-performing candidates with higher catenation levels compared to lower catenation levels.
Figure 8: Comparison of empirical cumulative distribution functions of the predicted CO2 capacities of generated and hMOF MOFs at different catenation levels.
Out of the 120,000 hypothetical MOF structures (see Section 4.3 for MOF assembly details), a total of 6,020 were predicted to be high-performing MOFs with CO2 capacity higher than 2 mmol/g at 0.1 bar. Categorization of the predicted high-performing MOFs by node-topology pair and catenation level (see Table 3) shows that the majority of Cu paddlewheel-pcu and Zn tetramer-pcu are cat2 and cat3, whereas the majority of high-performing Zn paddlewheel-pcu are cat1 and cat2. Table 4 shows the building blocks and catenation levels in the top five pcu MOF candidates. For linkers, the corresponding molecules (without dummy atoms) are shown for ease of visualization. Note that some substructures appear more often than others. For example, ring structures appear in most of the linkers in the top five candidates, and many of the DiffLinker-generated molecular substructures (in between the terminal fragments) also contain ring structures. The frequent occurrence of rings in linkers of high performing MOFs may be due to the presence of aromatic linkers, as reported by phenylene et al. [52].
We compared the chemical similarity of the linkers in predicted high-performing MOF structures with those in high-performing hMOF structures. As shown in Figure 13, we find that most linkers in the predicted high-performing MOF structures are vastly different from those in hMOF, indicating that the GHP-MOFassemble framework generate novel MOF structures with chemically unique linkers.
Table 3: Numbers of predicted high-performing MOFs found in the hMOF dataset.
node-topology cat0 cat1 cat2 cat3
Cu paddlewheel-pcu 0 170 1,184 628
Zn paddlewheel-pcu 9 789 929 395
Zn tetramer-pcu 0 3 794 1,119
Table 4: Details of the five pcu MOF candidates generated by GHP-MOFassemble with the highest predicted CO2 capacities. Their stability was confirmed using molecular dynamics simulations, and their 3D structures are shown in Figure 10.
5 Validation of AI-generated MOFs with molecular dynamics simulations
We now examine the stability and porous properties of the 6,020 MOFs described in the previous section by using molecular dynamics (MD) simulations. We use the LAMMPS [53] code, version 28 Mar 2023 Update 1, distributed from the conda-forge channel, to perform MD simulations that relax the AI-generated MOF structures. An adapted version of cif2LAMMPS [54] is used to generate LAMMPS input files automatically. For each MOF, a 2×2×2 supercell structure is generated, and an NPT simulation is performed, i.e., we use an isothermal-isobaric ensemble with a constant particle number $N$, a pressure $p$ fluctuating around an equilibrium value, $\langle p\rangle=1\textrm{atm}$, and a temperature $T$ fluctuating around an equilibrium value $\langle T\rangle=300\textrm{K}$. The cell lengths and angles of the triclinic simulation box are allowed to relax.
We run the NPT simulations for 200,000 steps with a timestep of 1 femtosecond. The relative changes of structure density before and after the simulation are inspected to evaluate how well the porous characteristics are maintained throughout the simulation. This process produces 102 MOFs whose structures stayed intact during the MD simulation. Next, we use a $<$1% change in density as the threshold to determine whether a MOF structure is stable or not. Among the 102 MOF candidates, 18 have change in density below the 1% threshold. Figure 9 presents the top five AI-generated, stable MOF structures that passed the threshold for structure determination—see Table 4 for combinations of their building blocks and Figure 10 for 3D visualization of their crystal structures.
Figure 9: Top five AI-generated MOFs with structure density changed less than 1% during molecular dynamics simulations. Their CO2 capacities are higher than 96.9% of hMOF structures in the hMOF database.
Figure 10: 3D visualization of the crystal structures of top five AI-generated, stable MOF candidates that retained their porous properties throughout molecular dynamics simulations. Atoms are presented using the color code: carbon in grey, nitrogen in dark blue, fluorine in cyan, zinc in purple, hydrogen in white, and lithium in green.
6 Conclusion
We have introduced GHP-MOFassemble, an AI-driven discovery framework for high-throughput generation of novel high-performing MOF structures for carbon capture. The GHP-MOFassemble framework employs a diffusion model, DiffLinker, to generate new, unique, and chemically diverse MOF linkers, which it then assembles with one of three pre-selected metal nodes into MOFs in the pcu topology.
Out of 120,000 MOF structures generated by GHP-MOFassemble, 6,020 were predicted to be high-performing MOFs with CO2 capacity higher than 2 mmol/g at 0.1 bar, which corresponds to the top 5% of the hMOF dataset. We relaxed each of these 6,020 MOFs in a 200-picosecond NPT molecular dynamics simulation with the UFF4MOF force field [55, 56], and found that 102 structures did not collapse after molecular dynamics simulations, among which 18 were identified to be stable structures with less than 1% change in density.
We have deployed and extensively tested GHP-MOFassemble on computers at the Argonne Leadership Computing Facility and at the National Center for Supercomputing Applications, with the intent of providing scalable and computationally efficient AI tools to accelerate the modeling and discovery of novel MOF structures. The tools introduced in this work may be readily fine-tuned and adapted to other available datasets beyond hMOF to enable accelerated design and discovery of novel MOF structures for carbon capture at industrial scale.
7 Acknowledgments
Code availability
The code and data are available upon reasonable request from the authors.
ORCID IDs
Hyun Park $0$-0001-5550-5610
Xiaoli Yan $0$-0002-6512-2338
Ruijie Zhu $0$-0001-9316-7245
Eliu Huerta $0$-0002-9682-3604
Santanu Chaudhuri $0$-0002-4328-2947
Ian Foster $0$-0003-2129-5269
Emad Tajkhorshid $0$-0001-8434-1010
Appendix A Dummy atom identification
Figure 11 illustrates the workflow used to identify the dummy atoms and to remove redundant atoms for linkers containing carboxyl groups (left panel) or heterocyclic rings (right panel).
Figure 11: Left panel, identification of dummy atoms for linkers containing carboxyl groups. The dummy atoms are found by substituting the carbon atoms in the carboxyl groups. The remaining oxygen and hydrogen atoms in the carboxyl groups are removed. Right panel, identification of dummy atoms for linkers containing heterocyclic rings. The dummy atoms are found at nitrogen-metal bond distance from the terminal nitrogen atoms along the vectors pointing from the opposing carbon atoms to nitrogen atoms.
Appendix B MOF catenation via site translation
Catenated MOFs were generated by using the site translation method, which is achieved by displacing the reference lattice along the diagonal line of the unit cell. For all four catenation levels, the amount of relative lattice displacement is given in Table 5. The numbers are fractional displacements relative to the unit cell diagonal.
Table 5: Amount of lattice displacement for catenated MOFs
catenation lattice1 lattice2 lattice3 lattice4
cat0 (0,0,0) / / /
cat1 (0,0,0) ($\frac{1}{2}$,$\frac{1}{2}$,$\frac{1}{2}$) / /
Appendix C Linker evaluation
@@
Recall from Section 4.2 that we use the synthetic accessible score (SAscore) and synthetic complexity score (SCscore) metrics to confirm that all linkers are synthesizable. Table 2 provides additional data on the generated linkers. We also provide results for the validity and uniqueness metrics, and the internal diversity scores IntDiv1 and IntDiv2 [48]. The internal diversity scores were calculated based on the Tanimoto distances among all pairs of molecules, which are obtained by calculating the normalized Jaccard score of Morgan fingerprint bit vector between all pairs of generated linkers. This yields similarities between a specific linker and the rest of the linkers in the linker pool, which are averaged out. Therefore, we end up with a metric showing each linker’s similarity compared with the population of generated linkers. We used the MOSES [48] framework to compute these two scores using the relations:
@@
where $G$ is the generated set of molecules, $|G|$ is the size of molecule set, $(m_{1},m_{2})$ represents a pair of molecules in the molecule set. $T_{d}$ is the Tanimoto distance, which relates to the Tanimoto similarity ($T_{s}$) by:
$T_{d}(m_{1},m_{2})=1-T_{s}(m_{1},m_{2})$ (9)
Appendix D Statistics and distribution of the standard deviation of the ensemble of pre-trained models
Table 6 shows the R2 score, mean absolute error (MAE), and root mean squared error (RMSE) of the three pre-trained models used for creating the ensemble model. Figure 12 shows the distribution of the standard deviations of ensemble model predictions.The threshold 0.2 mmol/g serves as a criterion for assessing whether the predictions made by the three pre-trained models are in agreement. The ensemble model predictions with standard deviation above the threshold were discarded.
Table 6: Statistics of the three pre-trained models
@@ Appendix D Statistics and distribution of the standard deviation of the ensemble of pre-trained models
Figure 12: Distribution of the standard deviation of ensemble model predictions. The ensemble model predictions with standard deviation of less than 0.2 mmol/g consist of around 96% of the test set. This shows that our three independently trained models are in great agreement for most of the unobserved test set, showing robustness of our model. The remaining 4% with larger than 0.2 mmol/g standard deviation implies that these are the data points which may be difficult to predict due to errors in target property or extra information other than atomic species and periodic neighbor is necessary for accurate prediction.
Appendix E Linker similarity for high-performing MOF structures
To measure the similarity between the generated linkers and hMOF linkers in high-performing MOF structures, we show in Figure 13 the distribution of maximum Tanimoto similarity between the two linker sets. For each unique linker in the predicted high-performing MOF structures generated by the GHP-MOFassemble framework, we calculate its Tanimono similarity with all of the unique linkers in the hMOF structures. The maximum value of the Tanimoto similarities gives a quantitative measure of how different the generated linkers are from the hMOF linkers.
Figure 13: Distribution of maximum Tanimono similarity between generated and hMOF linkers in high-performing MOF structures. The peak around 0.3 to 0.4 indicates that most generated linkers are just 30–40% similar to those in hMOF–showing that we are generating novel linkers not present in hMOF structures. On the other hand, the trailing heavy right tail above 0.4 Tanimoto similarity indicates that we are also able to generate linkers that are structurally similar to those present in hMOF, showing that GHP-MOFassemble enables generation of a diverse set of novel linkers.
Appendix F Regression model training details
We independently trained three modified versions of the CGCNN model for 5,000 epochs with a batch size of 160, and with optimizer, learning rate, and weight decay of torch_adam, 1e-4, and 2e-5, respectively.
Appendix G Linker statistics
Table 7: Statistical summary of the number of linkers after each step categorized by the corresponding MOF types. PW and TM stand for paddlewheel and tetramer, respectively. Element filter means that linkers with S, Br and I elements are removed.
Method Total Cu PW-pcu Zn PW-pcu Zn TM-pcu
DiffLinker 64,800 21,600 19,440 23,760
Hydrogen addition 56,257 18,979 17,126 20,152
Dummy atom identification 16,162 4,964 4,450 6,748
Element filter 12,305 3,702 3,441 5,162

### v2

@@ Methods
0.994
0.750 0.737
@@
The internal diversity scores were calculated based on the Tanimoto distances [62] among all pairs of molecules, which are obtained by calculating the normalized Jaccard score of Morgan fingerprint bit vector between all pairs of AI-generated linkers. This yields similarities between a specific linker and the rest of the linkers in the linker pool, which are averaged out. Therefore, we end up with a metric showing each linker’s similarity compared with the population of generated linkers. We used the MOSES [61] framework to compute the internal diversity scores IntDiv1 and IntDiv2 [61] by using the relations:
@@
where $T_{d}$ is the Tanimoto distance, which relates to the Tanimoto similarity ($T_{s}$) by:
$T_{d}(m_{1},m_{2})=1-T_{s}(m_{1},m_{2}),$ (9)
where $G$ is the generated set of molecules, $|G|$ is the size of that set, and $(m_{1},m_{2})$ is a pair of molecules in the set.
To measure the similarity between the AI-generated linkers and hMOF linkers in high-performing MOF structures, we show in Figure 10 the distribution of maximum Tanimoto similarity between the two linker sets. For each unique linker in the 364 AI-generated high-performing MOF structures we calculated its Tanimono similarity with all of the unique linkers in the hMOF structures. The maximum value of the Tanimoto similarities gives a quantitative measure of how different the generated linkers are from the hMOF linkers. Results in Figure 10 indicate that most linkers in the predicted high-performing MOF structures are vastly different from those in hMOF, i.e., our GHP-MOFassemble framework generates novel MOF structures with chemically unique linkers.
Figure 10: Similarity between AI-generated and hMOF linkers in high-performing MOF structures. Distribution of maximum Tanimono similarity, which measures the uniqueness of AI-generated linkers as compared to hMOF linkers for high-performing MOFs. The peak around 0.3 to 0.4 indicates that most generated linkers are just 30–40% similar to those in hMOF, i.e., our AI framework generates novel linkers not present in hMOF structures. On the other hand, the trailing heavy right tail above 0.4 Tanimoto similarity indicates that we are also able to generate linkers that are structurally similar to those present in hMOF, showing that GHP-MOFassemble enables generation of a diverse set of novel linkers.
MOF Assembly
Once we have generated linkers, we can then assemble them with metal nodes. To construct MOFs, we need to guide the assembly process. We do this by using dummy atoms, which indicate the points at which the building blocks are to be connected. In practice, this is done as follows. Our parsing of MOFids generates, for MOFs with Zn tetramer nodes, linkers that carry two carboxyl groups, which by definition are part of metal nodes instead of linkers. Such wrong assignment of carboxyl groups is due to how the MOFid algorithm parses MOF structures. To reflect the correct molecular structure of linkers, the two carbon atoms in the carboxyl groups are identified as dummy atoms, and the redundant four oxygen atoms and two hydrogen atoms are removed. An illustration of dummy atom identification for linkers containing carboxyl groups is shown in the left panel of Figure 11.
Figure 11: Assembly of AI-generated linkers with metal nodes. a Identification of dummy atoms for linkers containing carboxyl groups. The dummy atoms are found by substituting the carbon atoms in the carboxyl groups. The remaining oxygen and hydrogen atoms in the carboxyl groups are removed. b Identification of dummy atoms for linkers containing heterocyclic rings. The dummy atoms are found at nitrogen-metal bond distance from the terminal nitrogen atoms along the vectors pointing from the opposing carbon atoms to nitrogen atoms.
More than half of the hMOF structures are catenated MOFs, i.e., MOFs with interpenetrated lattices. By varying the level of interpenetration, it is possible to generate MOFs with different pore sizes, with a higher catenation level generally corresponding to smaller pores. We denote the four catenation levels in hMOF, with increasing number of interpenetrated lattices, as cat0, cat1, cat2 and cat3. To generate MOFs with high structural diversity, we applied site translation method as implemented in Pymatgen [63] to generate MOFs with different catenation levels. Figure 12 presents the building block combinations of the final six MOF candidates that passed all screening processes. For linkers, the corresponding molecules (without dummy atoms) are shown for ease of visualization. For catenated MOF structures, we keep the spacing between interpenetrated lattices the same, so as to ensure equal pore sizes.
Figure 12: Optimal linkers and catenation levels for MOF assembly. Building blocks and catenation levels of the top six AI-generated MOF candidates.
Screening and predicting properties of AI-generated assembled MOFs
Here we describe the final Screen and Predict component of GHP-MOFassemble. After the MOF structures are assembled, a comprehensive screening workflow is developed to sift through the candidates with increasing levels of computational cost and confidence level:
  • Inter-atomic Distance Check (screen)
  • Pre-simulation Check (screen)
[…]
  • Structure Validation of the Assembled MOFs (screen)
  • Grand Canonical Monte Carlo Simulation (screen and predict)
Inter-atomic Distance Check
A pairwise distance matrix $M_{n\times n}$ is extracted from a MOF’s CIF file. The matrix element $M_{i,j}$ denoting the Euclidean distance between atom $A_{i}$ and atom $A_{j}$ is:
$M_{i,j}=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}+(z_{i}-z_{j})^{2}}\,,$ (10)
where $1\leq i\leq n$ and $1\leq j\leq n$, and $(x_{i},y_{i},z_{i})$ and $(x_{j},y_{j},z_{j})$ are the Cartesian coordinates of atoms $A_{i}$ and $A_{j}$. For an arbitrary pair of atoms $A_{i}$ and $A_{j}$, their chemical symbols can be denoted as $E_{i}$ and $E_{j}$, respectively. The bond length between atoms $A_{i}$ and $A_{j}$ is $\delta(E_{i},E_{j})$. The minimum allowed inter-atomic distance between the these two atoms can be estimated by calculating the infimum of all experimental bond lengths between these two element types, i.e.: $\inf{(\delta(E_{i},E_{j}))}$, which can be queried from the OChemDb [64] database. All values of $M_{i,j}\ (i\neq j)$ are compared against $\inf{(\delta(E_{i},E_{j}))}$, and any occurrence of $M_{i,j}<\inf{(\delta(E_{i},E_{j}))}\ (i\neq j)$ will result in the MOF candidate being discarded. Entries with $i=j$ (i.e., the diagonal entries of the distance matrix $M_{i,i}$) are not examined as they should always be zero, because the distance between an atom and itself is zero. We used Python 3.10’s built-in multiprocessing library to help launch the inter-atomic distance check on different MOF structures in parallel.
Pre-simulation Check
For each atom in the MOF structure, the pre-simulation check workflow can be described as the following:
  1. 1.
    The atom’s element should be recognized as one of the supported elements in the UFF4MOF parameter set,
  2. 2.
    The atom’s neighboring atoms that are within the allowed bonding distance are examined for chemistry validity.
We conducted this process with the help of the cif2lammps library. Python’s multiprocessing library is also used to speed up this process over many structures in parallel.
Predicting $\textrm{CO}_{2}$ capacity of MOF via pre-trained regression model
The GHP-MOFassemble framework uses an ensemble of regression models to estimate the CO2 capacity of newly generated pcu MOF structures. We use for this purpose a modified version of the CGCNN model, developed in our previous work [37], which adopts an adjacency list to format node and edge embeddings, rather than the adjacency matrix format of the original CGCNN, a change that enhances model training speed, training stability, and prediction accuracy. To fine-tune this AI model, we used MOF structures in the hMOF dataset, along with their CO2 capacities at 0.1 bar as input. We split the hMOF dataset into 80% training set, 10% validation set, and 10% test test. We independently trained three modified versions of the CGCNN model for 5,000 epochs with a batch size of 160, using Adam as the optimizer, at a learning rate of $10^{-4}$, and a weight decay rate of $2\times 10^{-5}$.
Table 5 shows the R2 score, mean absolute error (MAE), and root mean squared error (RMSE) of the three AI models ensemble on the 10% test set. Figure 13 shows the distribution of the standard deviations of ensemble model predictions. We treat the standard deviation of predictions from the three models as a measure of model uncertainty, which we view as arising from the model’s difficulty in learning certain data points. This uncertainty is also known as epistemic uncertainty. By employing an ensemble of models and averaging their prediction results, we can increase prediction accuracy, which is thoroughly and extensively explored in the machine learning community [65, 66]. We chose the threshold for the standard deviation filter to be 0.2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ because it is sufficiently small (with 96% of data points below the threshold) that for low CO2 capacity predictions, the predicted values across the ensemble are similar. On the other hand, for high CO2 capacity predictions, it also ensures that the outliers (i.e., predicted values across the ensemble are vastly different) are filtered out, therefore minimizing the overall error. In Table 6 we also notice that most of the high-performing MOFs are cat2 and cat3. This is consistent with our preliminary analysis on catenation levels and distributions of CO2 adsorption capacity in Figure 5.
Table 5: Statistical properties of AI ensemble to characterize MOF’s performance. Statistics of AI ensemble, including R2 score, mean absolute error (MAE), and root mean squared error (RMSE).
@@
Figure 13: Standard deviation of AI model ensemble to estimate MOFs’ CO2 capacity. Distribution of the standard deviation of AI ensemble predictions. Predictions with standard deviation of less than 0.2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ consist of around 96% of the test set. This shows that our three independently trained AI models are in great agreement. The remaining 4% with larger than 0.2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ standard deviation implies that these are the data points which may be difficult to predict due to errors in target property or extra information other than atomic species and periodic neighbor being necessary for accurate predictions.
Table 6: Catenation level of high-performing MOFs. Numbers of predicted high-performing MOFs found in the hMOF dataset.
node-topology cat0 cat1 cat2 cat3
Cu paddlewheel-pcu 0 1 38 26
Zn paddlewheel-pcu 0 0 66 23
Zn tetramer-pcu 0 0 53 157
We show in the left panel of Figure 14 the scatter plot of the ensemble model predictions for the 10% test set (from hMOF), which has a MAE of 0.093 m $\textrm{m\,mol}\,\textrm{g}^{-1}$. Using CO2 capacity of 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ as a threshold, we repurposed our predictive model as a classifier to categorize MOFs into low and high performers, with predicted CO2 capacities below and above the threshold, respectively. Using this scheme, the confusion matrix for identifying low and high performers is shown in the right panel of Figure 14. We conclude from the confusion matrix that the pre-trained model classifies both low and high performers with high accuracy, with 98.4% (13,551 out of 13,765) of test samples (of hMOF) correctly classified. As the test set is heavily imbalanced, with many more low performers than high performers, we also calculated the balanced accuracy of classification [67] (the sum of true positive rate and true negative rate divided by two for binary classification or average of recalls for multi-class classification), obtaining a value of 90.7%. Since the majority of the 214 misclassified samples lie close to the decision boundaries, as shown in the red and purple regions of the left panel of Figure 14, we conclude that the ensemble model is capable of differentiating low and high performers.
Figure 14: Performance of AI ensemble to measure CO2 capacity. a Predictive performance of the ensemble model on the 10% test set of hMOF. b Confusion matrix of the ensemble model in classifying low and high performers in the test set. MOFs with real/predicted CO2 capacity lower than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ are identified as low performers, and MOFs with real/predicted CO2 capacity higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ are identified as high performers.
We show in Figure 15 the empirical cumulative distribution functions of the CO2 capacities of hMOF (dashed lines) structures and AI-generated MOF structures (solid lines) at different catenation levels. We observe that there are more high-performing cat2 and cat3 MOFs among the generated structures compared to the hMOF dataset, as indicated by lower cumulative density values at 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$ (black dashed line). On the other hand, there are fewer high-performing cat0 and cat1 MOFs, as shown by higher cumulative density values at 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$. This result demonstrates that within the AI-generated MOF design space, there are more high-performing candidates with higher catenation levels compared to lower catenation levels.
Figure 15: MOFs’ CO2 capacities in terms of catenation levels. Comparison of empirical cumulative distribution functions of the predicted CO2 capacities of generated and hMOF structures for a cat0, b cat 1, c cat2, and d cat 3.
In summary, we use an ensemble of 3 AI models to identify pcu AI-generated MOFs produced by the Generate step (in addition to inter-atomic distance and pre-simulation checks) that have predicted CO2 capacities higher than 2 m $\textrm{m\,mol}\,\textrm{g}^{-1}$. We use the ensemble mean plus standard deviation as the final prediction. These MOFs are then selected for further investigation with MD simulations.
Structure Validation of the Assembled MOFs
Pre-screened MOFs are checked for stability using MD simulations with LAMMPS [42], version stable release 2 August 2023 compiled with gcc 11.2.0 and OpenMPI 4.1.2. For each MOF, a 2x2x2 supercell of triclinic periodic boundary condition is created. A triclinic isothermal-isobaric ensemble (i.e., NPT) at 300 K and 1 atm is applied to the supercell structure. All lattice parameters: $a$, $b$, $c$, $\alpha$, $\beta$, $\gamma$ are allowed to equilibrate freely. A relatively short simulation with 400,000 steps is conducted for each MOF to allow for structural equilibrium with a step size of 0.5 femtoseconds. The potential energy $E_{\textrm{pot}}$ of the MOF structure can be expressed as:
$E_{\textrm{pot}}=E_{\textrm{bond}}+E_{\textrm{angle}}+E_{\textrm{dihedral}}+E_{\textrm{improper}}+E_{\textrm{Vdw}}+E_{\textrm{Coul}},$ (11)
where $E_{\textrm{bond}}$ is the bond stretch energy, $E_{\textrm{angle}}$ is the angle between bonds energy, $E_{\textrm{dihedral}}$ is the dihedral angle energy, $E_{\textrm{improper}}$ is the improper dihedral angle energy, $E_{\textrm{vdw}}$ is the nonbonded interaction energy, and $E_{\textrm{coul}}$ is the Coulombic interaction energy.
Coupry et al. (2016) [40] stated that 76.5% of the CoRE structures [14] lattice parameters are within 5% change when they are simulated with the UFF4MOF parameters. A similar criterion is used here such that if any of MOF’s lattice parameters of triclinic cell changes are more than 5%, the structure is discarded in the screening process. This screening process provides a new level of confidence in terms of structural stability of the AI-generated MOFs. Figure 16 demonstrates the relative error of all lattice parameters of the top six MOF candidates during the MD simulation.
Figure 16: Lattice parameter relative errors of the top 6 AI-generated MOFs. The top six AI-generated MOF candidates have change less than 5% in all lattice parameter during MD simulation. The first three lattice parameters are cell axes length and the latter three lattice parameters are cell plane angles. Their CO2 capacities are higher than 96.9% of MOF structures in the hMOF database.
Grand Canonical Monte Carlo (GCMC) Simulations
The void fraction of each candidate MOF is calculated using the RASPA [43]’s helium void fraction function. It helps us understand the porosity contribution toward MOF CO2 adsorption capacity. Next, PACMOF [68] is used to assign partial charges to MOFs. Using the pre-trained partial atomic charge model on the density derived electrostatic and chemical (DDEC) charge [69] data, the partial charges are assigned to atoms in MOF structures. After determining the partial charges, the MOF structures are assigned with UFF4MOF force field parameters. The GCMC simulations are conducted at 0.1 bar and 300 K with a step size of 0.5 fs/step. The MOF structures are under the rigid assumption, therefore only Van der Waals interactions and Coulombic interactions (i.e., non-bonded) are considered:
$E_{\textrm{pot}}=E_{\textrm{Vdw}}+E_{\textrm{Coul}},$ (12)
which can be expressed as [70]
$E_{\textrm{pot}}=\sum_{i,j}{4\epsilon_{ij}\bigg{[}\Big{(}\frac{\sigma_{ij}}{r_{ij}}\Big{)}^{12}+\Big{(}\frac{\sigma_{ij}}{r_{ij}}\Big{)}^{6}\bigg{]}}+\frac{q_{i}q_{j}}{4\pi\epsilon_{0}r_{ij}},$ (13)
where $\epsilon_{ij}$, $\sigma_{ij}$, and $r_{ij}$ are the well depth, zero position, and inter-atomic distance of a Lennard-Jones 12-6 interaction between atom $i$ and atom $j$; $q_{i}$ and $q_{j}$ are the electrostatic charges of atom $i$ and atom $j$; $\epsilon_{0}$ is the vacuum permittivity.
Computational resources and performance
We have deployed and extensively tested GHP-MOFassemble on computers at the ALCF and at the National Center for Supercomputing Applications (NCSA), with the intent of providing scalable and computationally efficient AI tools to accelerate the modeling and discovery of novel MOF structures. The tools introduced in this work may be readily fine-tuned and adapted to other available datasets beyond hMOF to enable accelerated design and discovery of novel MOF structures for carbon capture at industrial scale. The training and inference of the generative model and the training of the regression model was conducted on 8-way NVIDIA A100 GPUs with FP32 mode. The inference of the regression model is conducted on a single NVIDIA A40 GPU with FP32 mode. CPU-based screening processes, including distance check, pre-simulation check, MD simulations, and GCMC simulations, are run on two-way AMD Epyc 7763 CPUs. We now present a breakdown of the average computational cost of each step of the GHP-MOFassemble workflow per MOF:
  • MOF assembly: 0.46 seconds per core. We scaled up this analysis by multiprocessing on 28 cores.
  • Distance check to ensure structural validity: 2.56 seconds per core. We scaled up this analysis by multiprocessing on 128 cores.
  • Pre-simulation check to ensure chemical consistency: 19.98 seconds per core. We scaled up this analysis by multiprocessing on 128 cores.
  • AI ensemble inference of CO2 adsorption capacity: 20.46 seconds. We used one NVIDIA A40 GPU.
  • MD simulations to validate stability and chemical consistency: 658.40 seconds, or about $\sim 10$ minutes. We scaled up this analysis using between six to 14 MPI processes based on the number of atoms in the MOF.
  • Detailed GCMC simulations: 21214.04 seconds, or about $\sim 6$ hours. This averaged number is based on simulations on one CPU core.
A schematic representation of this benchmark analysis is presented in  Figure 17.
Figure 17: Process time cost breakdown of GHP-MOFassemble. Benchmark analysis of the computational cost for each process of the GHP-MOFassemble framework. Notice that the AI component is hyperefficient.
Data Availability
The datasets generated during and/or analysed during the current study are available in the GitHub repository, https://github.com/hyunp2/ghp_mof/tree/main/utils. We also used the open source hMOF dataset [71], and the GEOM dataset [56].
Acknowledgements (not compulsory)
Code availability
The scientific software and data used in this article are readily available in GitHub at https://github.com/hyunp2/ghp_mof/tree/main.
Author contributions statement
E.A.H. envisioned and led this work, and guided the connection between generative AI with MD and GCMC simulations to create novel, chemically consistent MOFs at scale in high performance computing platforms. H.P. adapted techniques from drug design into MOF discovery with generative AI, producing a pool of candidates of MOF components, and also trained and inferred assembled MOFs’ CO2 adsorption capabilities using a graph neural network model. X.Y. developed codes to decompose MOF into linkers and nodes, as well as to assemble predicted linkers and nodes to make hypothetical MOFs, and also conducted large-scale MD simulations, with LAMMPS and GCMC to identify stable and synthesizable AI-predicted MOFs. R.Z. analyzed hMOF database to understand MOF composition statistics, tested combined generative AI and graph models to predict novel AI-generated MOFs and quantify their chemical properties. S.C. provided expertise on metrics to be used to study the chemical properties and stability of MOFs with AI tools developed in this manuscript. D.C. provided expertise on expected features of MOFs to be used for carbon capture at industrial scale. I.F. guided the development of our ready-to-use AI framework in modern computing environments, and provided guidance on how to use and interpret metrics for MOF synthesizability. E.T. provided expertise on MD, chemistry and synthesis pathways of MOFs as well as how to analyze MD results. All authors reviewed and contributed to the manuscript.
Corresponding author
Correspondence to Eliu Huerta.
Ethics declarations
Competing interests
The authors declare no competing interests.